OpenStudy (anonymous):

What is the domain of y = tan(x) + 2

OpenStudy (anonymous):

domain is \[R - {\pi/2, 3\pi/2 ,...,(2k+1)\pi/2}\]

OpenStudy (anonymous):

numbers are in R- { }